OpenStudy (anonymous):

[8.04] Part 1: Tommy and Jessica are discussing how to factor 2x2 + 15x + 28. Tommy feels this trinomial is prime because he cannot find the factors of 28 that have a sum of 15. Jessica says he is incorrect and that it is factorable. Using complete sentences, provide a convincing argument explaining who is correct and why. If this trinomial is factorable, factor it showing all work and explain your steps. Part 2: Create your own prime trinomial in the form ax2 + bx + c. Using complete sentences, explain how you know it is prime. I just need help on part two

OpenStudy (anonymous):

help on this part Part 2: Create your own prime trinomial in the form ax2 + bx + c. Using complete sentences, explain how you know it is prime.

OpenStudy (austinl):

\(x^2-3x+15\) A way you can tell if it's prime or not is to feed it into a portion of the quadrac formula: \(\sqrt{b^2 - 4ac}\) If the result is an integer, it's not prime (aka it can be factored). If the result is something in radical form, then it is prime. So in this case: \(\sqrt{(-3)^2 - 4(1)(15)}\) \(\sqrt{6 - 60}\) \(\sqrt{-54}\) The result are imaginary numbers, so it cannot be factored, thus it's prime.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

is that the answer??????